dy/dx = y(20 - 4y)


1
Expert's answer
2021-03-02T00:51:10-0500
"\frac{dy}{dx}=y(20-4y)""\frac{dy}{y(5-y)}=4dx""\int\frac{dy}{y(5-y)}=\int4dx""\int\frac{1}{5}(\frac{1}{y}+\frac{1}{5-y})dy=\int4dx""\frac{1}{5}\ln |y|-\frac{1}{5}\ln|y-5|=4x+C,"

where "C" is some constant.



"\ln{\frac{y}{y-5}}=20x+C_1,"

or


"y=\frac{5}{C_2e^{-20x}+1}."

Answer. "y=\frac{5}{C_2e^{-20x}+1}."